# COBias and Debias: Minimizing Language Model Pairwise Accuracy Bias via Nonlinear Integer Programming

### v1

@@ 3.2 Solving DNIP with Simulated Annealing
We solve DNIP by simulated annealing [21, 22]. Simulated annealing (SA) is a Metropolis-Hastings sampling algorithm, which is widely adopted to solve optimization problems with its versatility in dealing with highly nonlinear models, noisy data, and many constraints [23]. The DNIP mathematical model is such a case, pointing to the need of SA. We especially take advantages of SA algorithms’ searching ability - escaping from local optima and finding global optima after reaching local optima - to find optimal correction weights. In details, the SA strategy consists of searching the solution space starting from a randomly initialized $\xi$ and then generating a new one by perturbing it. The new solution is either accepted or rejected by the Metropolis criterion after evaluating the objectives.
Algorithm 1 describes the searching for optimal $\xi$ using SA. The outer loop performs the cooling/annealing process. In this work, a geometric decay with $\alpha=0.95$ and initial temperature of 200,000 is set as cooling schedule. The inner loop simulates the thermal equilibrium reaching process at a temperature, and its criterion ensures that either the number of generated or accepted solutions at a temperature is above a threshold [24]. We sample a new $\xi$ from the neighborhood by randomly substitute a selection in $\xi$. The computational complexity for the SA algorithm is $\mathcal{O}(NK)$. Please refer to the Appendix Section A for derivations.
@@ 4.1 Experimental Setup
Models and Experimental Configuration. We evaluate three open-source LLMs across varied scales and model families, including GPT-2-XL (1.5B parameters), Llama-2-7B (7B parameters), and Llama-2-13B (13B parameters). For ICL, we use 1-shot prompting, and obtain the output per-class softmax probabilities. To standardize calculations, we normalize the softmax probabilities over all classes. For DNIP, we tune parameters of $\beta,\tau,K,u$ on the development set. To account for variance caused by different demonstrations, for each model and dataset, we perform three runs of prompting with different 1-shot demonstrations and obtain three sets of initial probabilities. For both ICL and DNIP, we report the average test accuracy and COBias over the three runs. The demonstration is randomly selected from training examples. All prompting is done on an NVIDIA A100 GPU. The simulated annealing algorithm executes on CPU (execution time depends on the optimization set size, 3 to 30 minutes on AMD EPYC 7742 CPU or slightly longer on local laptop CPU).
@@ 4.2 Main Results
The average test accuracy (in black) and COBias (in blue) with standard deviation over three runs are shown in Table 1. We find that test COBias by ICL is large in all three LLMs, with GPT-2-XL having the largest average COBias of 50.3%, showing that LLMs of different sizes and families exhibit large per-class accuracy differences. With DNIP, the average test COBias over three models reduces from 43% to 16%, and the test accuracy increases from 52% to 64%. Specifically, DNIP achieves an 16%, 12%, and 10% absolute improvement in test accuracy for GPT-2-XL, Llama-2-7B, and Llama-2-13B respectively; meanwhile, DNIP significantly reduces test COBias by an absolute 29%, 23%, and 26% for the three models, over ICL results. As for detailed results, for GPT-2-XL where the ICL test accuracy is below 30% (DDI), DNIP boosts test accuracy by 29% while lowering test COBias from 41% to 17%; for Llama-2 7B and 13B where the ICL test accuracy is already in higher 80s, e.g., DBpedia, DNIP can further improve the accuracy to 90s while reducing COBias. In addition, we observe a very large ICL test COBias at 82.4% for GPT-2-XL on RTE. After applying DNIP, the COBias greatly reduces to 7.1%. The 82.4% COBias on RTE suggests that the conventional ICL predictions mostly fall in a single class, failing to predict the other class. In fact, The RTE task seems too challenging that GPT-2-XL’s output probability of class True is much higher than that of class False. Though re-weighting can lower both probabilities, it may not gain a corrected prediction. Still, the over 75% absolute reduction in COBias demonstrates that DNIP in the worst case can balance severely skewed ICL predictions without hurting the overall accuracy much. On two biomedical tasks, DNIP brings an average absolute increase of 34% and 11% in accuracy, and an average absolute reduction of 31% and 13% in COBias, showing its effectiveness for domain-specific tasks.
These results show that DNIP strongly mitigates COBias and greatly improves the accuracy of lower-performing classes, while boosting overall accuracy. Moreover, relatively large LLMs can suffer even higher COBias than smaller models on biomedical tasks, suggesting that the per-class accuracy bias does not go away as models scale, and debiasing is essentially needed to improve inference for larger LLMs.
@@
4.3 DNIP Balances Class Accuracies across Varied Classification Tasks, as many as 14 Classes
@@
Figure 2: Comparisons of pairwise class accuracy differences on test sets before and after applying DNIP; the pinker the higher difference, the greener the lower difference; average accuracy over three runs are used. Heatmaps after debiasing show nearly 0 differences for many class pairs, demonstrating DNIP’s effectiveness for COBias reduction.
We visualize pairwise class accuracy differences before and after applying DNIP on Llama-2-13B by the heatmaps in Figure 2, which show that DNIP is straightforwardly effective in reducing COBias. In each heatmap, the value of row $i$, column $j$ cell represents the absolute difference between the test accuracies of two classes $i,j$ , i.e., $|A_{i}-A_{j}|$. We observe that DNIP brings the accuracy differences down to close to zero for 4 out of 7 datasets, including AGNews, DDI, DBpedia, and RTE; and the most relative reduction is seen on TREC. These results show that DNIP consistently reduces prediction bias across different NLP classification tasks.
@@
Figure 3: Test confusion matrix for DBpedia before and after applying DNIP. DNIP greatly reduces wrong predictions for Nat., which is the lowest-accuracy class before debiasing.
@@
Relieving Odd Classes for 14-Way Ontology Classification. We further investigate the 14-way classification task DBpedia. Before debiasing, the Nat. class exhibits the largest class accuracy difference compared to other classes, suggesting that Nat. is either the highest-accuracy class manifesting a strong prediction bias towards Nat., or the lowest-accuracy class where predictions of true Nat. examples are biased to one or more odd classes. We plot the confusion matrix of test results of 1 run out of 3 runs in Figure 3 and find that Nat. has the lowest accuracy, and its predictions are biased to two odd classes, Trans. and Bldg., with Trans. being the oddest to Nat.. DNIP greatly reduces false negatives for Nat., i.e., reduces wrong Trans. and Bldg. predictions for Nat.. Besides these visualizations, on an average over three runs, the test accuracy for Nat. improves from 44% to 89%, and that for Trans. also increases from 78% to 87%, demonstrating DNIP’s effectiveness in reducing pairwise class accuracy bias for classification tasks of as many as 14 classes.
Furthermore, we see more balanced class accuracies with DNIP for smaller LLMs of Llama-2-7B and GPT-2-XL, in Appendix Section B.
4.4 DNIP Objective Ablation
@@
We present ablation analysis on objective functions to show that the integration of accuracy-related objectives and the COBias objective are indispensable in achieving the best of both worlds. Recall three parts in our objective function $z$: the error rate term $z_{1}=\frac{1}{M}\sum_{m=1}^{M}\mathbbm{1}\{\hat{y}_{m}\neq y_{m}\}$; the COBias term $z_{2}={N\choose 2}^{-1}\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}|A_{i}-A_{j}|$; and the PMI term $z_{3}=\sum_{j=1}^{N}\textrm{PMI}_{j}$; the accuracy-related terms are $z_{1}$ and $z_{3}$. We set up 7 different objective functions to perform DNIP: $z_{1},z_{2},z_{3},z_{1}+\beta z_{2},z_{1}-\tau z_{3},\beta z_{2}-\tau z_{3},z_{1}+\beta z_{2}-\tau z_{3}$, and compare their effectiveness on Llama-2-13B in Figure 4. Ablations on GPT-2-XL and Llama-2-7B show similar findings (Appendix Section C). To make fair comparisons, we fix $\beta=2.7,\tau=0.2$ and weight scale $K=30$, which may not result in the same scores reported in Table 1.
@@
Among all ablations, $z_{1}+\beta z_{2}-\tau z_{3}$ achieves a balance point between accuracy and COBias. For the criteria, lower COBias is our top priority; when COBias scores are similarly low, we prefer the one that achieves higher accuracy. Compared to $z_{1}+\beta z_{2}-\tau z_{3}$$z_{1}$: (1) Using $z_{1}$ or $z_{3}$ solely emphasizes improving accuracy; though obtaining higher or similar accuracy, they suffer a higher COBias. (2) Using $z_{1}-\tau z_{3}$ also obtains a higher/similar accuracy but a higher COBias. (3) Using $z_{2}$ solely focuses on mitigating COBias and is straightforwardly effective in reducing COBias. However, the shortcoming is obvious: a lower accuracy. (4) Using $z_{1}+\beta z_{2}$ or $\beta z_{2}-\tau z_{3}$ is promising with a similar/lower accuracy and a slightly lower COBias. We note that accuracy-related objectives alone increase accuracy by over 16% ($75\%-59\%$), which is greater than using the objective combining COBias, suggesting the NIP can be a direction for those who look for accuracy improvement.
@@
Test accuracy is improved with only 10 optimization examples, and plateaus at around 500 examples. Figure 6 shows that, at 10 examples, the relative increase in accuracy ranges from 0.2% to 11% except DDI, which is 160%. At 100 examples, the increase ranges from 3% to 208% (DDI, $7.2\%\rightarrow 18.7\%$). The increase continues to grow and stabilizes at around 500 examples. At 1,000 examples, test accuracy for DDI rises to 52%, which is more than 6 times of the ICL accuracy, showing that DNIP is effective for this challenging biomedical NLP task. Moreover, we observe for RTE a declining trend of accuracy increase, which matches its growing trend of COBias reduction, reflecting the trade-off between the two objectives. We also note a slight decrease in accuracy improvement for SST-5 and DDI at full size, which matches their sudden increase in COBias reduction, verifying that DNIP has emergent COBias reduction abilities with thousands of optimization examples. Overall, DNIP improves both accuracy and COBias at 10 examples, and achieves acceptable COBias reduction and accuracy improvement at 100 examples, with test COBias reduced by 57% and test accuracy improved. For detailed scores, please refer to Appendix D.
@@ 5 Related Work
LLMs may fail in various ways. Analysis towards understanding LLM failure modes has gained remarkable insights on how they fail and how to identify/avoid failures. For example, understanding hallucinations in LLMs facilitates fact-checking evaluation metrics for LLMs.[32, 33, 34]. Jones and Steinhardt [14] reveals failure modes of LLMs which resemble cognitive biases and measures the failures on GPT-3.5 and Codex. Lu et al. [9] identify ICL order sensitivity that result in LLM prompting failures. Turpin et al. [10] shows that LLM reasoning abilities are affected by adding biasing features to the input. Along this line, we tackle a common LLM failure of the per-class accuracy difference, via nonlinear integer programming optimization on the evaluation metric scores.

### v3

@@ 3.2 Solving DNIP with Simulated Annealing
We solve DNIP by simulated annealing [6, 7]. Simulated annealing (SA) is a Metropolis-Hastings sampling algorithm, which is widely adopted to solve optimization problems with its versatility in dealing with highly nonlinear models, noisy data, and many constraints [8]. The DNIP mathematical model is such a case, pointing to the need of SA. We especially take advantages of SA algorithms’ searching ability - escaping from local optima and finding global optima after reaching local optima - to find optimal correction weights.
In details, the SA strategy consists of searching the solution space starting from a randomly initialized $\xi$ and then generating a new one by perturbing it. The new solution is either accepted or rejected by the Metropolis criterion after evaluating the objectives.
Algorithm 1 describes the searching for optimal $\xi$ using SA. The outer loop performs the cooling/annealing process. In this work, a geometric decay with $\alpha=0.95$ and initial temperature of 200,000 is set as cooling schedule. The inner loop simulates the thermal equilibrium reaching process at a temperature, and its criterion ensures that either the number of generated or accepted solutions at a temperature is above a threshold [9]. We sample a new $\xi$ from the neighborhood by randomly substitute a selection in $\xi$. The computational complexity for the SA algorithm is $\mathcal{O}(NK)$. Please refer to Appendix C for derivations.
@@ 4.1 Experimental Setup
Models and Experimental Configuration. We evaluate three open-source LLMs across varied scales and model families, including GPT-2-XL (1.5B parameters), Llama-2-7B (7B parameters), and Llama-2-13B (13B parameters). For main results, we use 1-shot prompting and obtain the ICL output class softmax probabilities. To standardize calculations, we normalize the softmax probabilities over all classes. For DNIP, we tune parameters of $\beta,\tau,K,u$ on the development set. To account for variance caused by different demonstrations, for each model and dataset, we perform three runs of prompting with different 1-shot demonstrations and obtain three sets of initial probabilities. For both ICL and DNIP, we report the average test accuracy and COBias over the three runs. The demonstration is randomly selected from training examples. All prompting is done on an NVIDIA A100 GPU. Simulated annealing executes on CPU (execution time depends on the optimization set size, 3 to 30 minutes on AMD EPYC 7742 CPU or slightly longer on local laptop CPU).
@@ 4.2 Main Results
The average test accuracy (in black) and COBias (in blue) with standard deviation over three runs are shown in Table 3. In all three LLMs, test COBias by ICL is large, with GPT-2-XL exhibiting the largest COBias (50.3%), showing that class accuracy imbalance exists in LLMs of different sizes and families. With DNIP, test COBias is significantly reduced, by an absolute 29%, 23%, and 26% for the three models respectively; meanwhile, there is an 16%, 12%, and 10% absolute improvement in test accuracy respectively. It is worth noting that for tasks whose ICL test accuracy is already in higher 80s, e.g., DBpedia, DNIP can reduce COBias while further improving the accuracy to 90s. In addition, we observe a very large ICL COBias at 82.4% in GPT-2-XL for RTE, reflecting severe over-predictions in class True. After applying DNIP, the COBias greatly reduces to 7.1%. That is, DNIP obtains over 75% absolute reduction in COBias for RTE, demonstrating DNIP’s capability for balancing very imbalanced class accuracies without hurting the overall accuracy much. Moreover, on the two biomedical tasks, DNIP brings an average absolute reduction of 31% and 13% in COBias, and an average absolute increase of 34% and 11% in accuracy, showing its effectiveness for domain-specific tasks.
Moreover, the 13B Llama model can suffer even higher COBias than its 7B counterpart on biomedical tasks, further showing that class accuracy imbalance does not go away as model sizes increase, and DNIP is effective to balance class accuracies for LLMs of all sizes.
Back to the AGNews example in Table 2, we present the DNIP debiased results in Table 2. Over-predictions in Business are greatly reduced, and the accuracy of Tech is significantly improved from 19% to 84%. These results show that DNIP strongly mitigates COBias and greatly improves the accuracy of lower-performing classes, while boosting overall accuracy.
Computational Time. Regarding computational time needed to perform DNIP optimization, on all datasets, the optimization is finished in the scale of minutes. Though computational time increases with more classes and a more fine-grained weight scale, it should be noted that the time elapsed in optimization is dozens of minutes even for tasks of more than 10 classes, and a weight scale of 30 is good enough for most of the datasets experimented in this paper. This demonstrates the practicality of DNIP in the computational time needed for performing optimization.
Prompting Method Metric AGNews DBpedia SST-5 TREC RTE DDI PubMedQA Avg.
Acc $82.5_{2.0}$ $93.6_{1.3}$ $45.8_{6.0}$ $58.7_{23.3}$ $61.9_{16.9}$ $34.1_{42.1}$ $70.4_{7.1}$ 63.9
5-shot ICL COBias $24.4_{4.2}$ $9.0_{2.0}$ $48.0_{15.4}$ $35.4_{13.7}$ $61.3_{40.4}$ $44.4_{5.4}$ $52.2_{19.3}$ 39.2
Acc $88.5_{0.6}$ $95.8_{0.7}$ $52.9_{2.5}$ $76.6_{8.7}$ $75.8_{4.5}$ $54.1_{5.1}$ $59.9_{1.4}$ 71.9
5-shot DNIP COBias $7.0_{0.9}$ $5.7_{1.1}$ $15.8_{12.6}$ $14.4_{7.5}$ $3.1_{1.9}$ $16.5_{7.6}$ $9.6_{4.6}$ 10.3
Acc $83.5_{1.5}$ $95.2_{1.2}$ $50.3_{2.3}$ $67.0_{12.7}$ $75.0_{0.8}$ $9.7_{1.0}$ $52.3_{5.3}$ 61.9
k-shot (1 demon. per class) ICL COBias $14.9_{5.1}$ $7.0_{2.2}$ $36.3_{7.2}$ $38.2_{5.1}$ $22.5_{13.2}$ $39.7_{3.5}$ $20.9_{4.2}$ 25.6
Acc $88.7_{0.5}$ $96.6_{0.5}$ $51.3_{1.0}$ $82.7_{1.4}$ $76.7_{3.7}$ $43.6_{6.1}$ $55.3_{2.6}$ 70.7
k-shot (1 demon. per class) DNIP COBias $7.3_{0.5}$ $4.3_{0.7}$ $2.8_{1.5}$ $12.1_{5.5}$ $5.0_{3.3}$ $12.5_{4.3}$ $8.7_{1.2}$ 7.5
Table 4: DNIP results in more ICL settings. Average score over three runs are reported.
4.3 More ICL Settings
We prompted Llama-2-13B in two additional settings: 5-shot, where each demonstration is randomly selected as in the 1-shot case; k-shot, where k is the number of classes and each class is represented by a demonstration example in the prompt. Table 4 presents results on seven benchmark datasets in both settings. For example, the relative average COBias reduction (comparing DNIP to ICL) is 65% and 74% for 1-shot and 5-shot cases respectively, and the relative average accuracy increase is 16% and 11% for 1-shot and 5-shot cases respectively, demonstrating DNIP’s capabilities with more shots. In addition, prompt design does contribute to different output probabilities, which could be helpful to provide different starting points for DNIP, so DNIP could optimize for better solutions. However, prompt engineering alone may not be most effective to solve the COBias issue. For example, compared to 1-shot prompting (Table 3), increasing the number of shots does not significantly further reduce COBias or boost accuracy; using more diverse demonstrations (k-shot) may help with COBias, but does not gain much higher accuracy than 1-shot. Therefore, using more shots does not help reduce COBias much, and DNIP significantly reduces COBias and improves accuracy in both 5-shot and k-shot settings, further showcasing its effectiveness.
@@ 4.3 More ICL Settings
4.4 DNIP Significantly Reduces Over-predictions, and Balances Class Accuracies across Varied Classification Tasks, as many as 14 Classes
@@
Figure 2: Pairwise class accuracy differences on test sets before and after applying DNIP; the pinker the higher class accuracy difference, the greener the lower; average accuracy over three runs are used. Heatmaps after DNIP show nearly 0 accuracy differences for many class pairs.
@@
Figure 3: Test confusion matrix for DBpedia before and after applying DNIP. DNIP greatly improves ICL predictions for Nat., which is the most under-predicted class by ICL.
We visualize pairwise class accuracy differences before and after applying DNIP to 1-shot, Llama-2-13B ICL results in Figure 2. In each heatmap, row $i$, column $j$ cell represents the absolute difference between the test accuracies of two classes in row $i,j$ , i.e., $|A_{i}-A_{j}|$. DNIP brings the accuracy differences down to close to zero for AGNews, DDI, DBpedia, and RTE. The lighter green heatmaps in the bottom row show that DNIP consistently reduces COBias across different NLP classification tasks.
@@
Relieving Odd Classes for 14-Way Ontology Classification. We further investigate the 14-way classification task DBpedia. Before debiasing, the Nat. class exhibits the largest class accuracy difference to other classes. In fact, Nat. is the most under-predicted class, whose odd class is Trans. and second odd class is Bldg., as shown by the confusion matrix in Figure 3 (based on seed 0, 1-shot ICL run). DNIP greatly reduces false negatives for Nat., i.e., reduces wrong Trans. and Bldg. predictions for Nat.. Moreover, on an average over all three runs, the test accuracy for Nat. improves from 44% to 89%, and that for Trans. also increases from 78% to 87%, demonstrating that DNIP is effective in reducing class accuracy differences for tasks of as many as 14 classes. Furthermore, we present more heatmaps for smaller LLMs including Llama-2-7B and GPT-2-XL in Appendix D.
4.5 Optimizing COBiassingle
Using COBiassingle (Eq. 4) for DNIP optimization will focus on odd classes only. Figure 4 shows ablation results when replacing the COBias term in the DNIP objective with COBiassingle.
$\min\quad Z(\xi)=\frac{1}{M}\displaystyle\sum_{m=1}^{M}\mathbbm{1}\{\hat{y}_{m}\neq y_{m}\}+\beta\frac{1}{N}\displaystyle\sum\nolimits_{i}\bigg{|}A_{\textrm{odd}_{i}}-A_{c_{i}}\bigg{|}-\tau\displaystyle\sum_{j=1}^{N}\textrm{PMI}_{j}$ (14)
Figure 4: Test COBiassingle, COBias, and overall accuracy scores (from left to right).
Blue dots represent the respective evaluation metric score of the original 1-shot ICL Llama-2-13B outputs of the seven benchmark datasets; green dots are debiased results optimized by the COBiassingle DNIP objective (Eq. 14); red dots are debiased results optimized by the COBias DNIP objective.
Comparing DNIP with COBiassingle and DNIP with COBias, their performances are close, but DNIP with COBias achieves higher overall test accuracy and slightly better bias mitigation performances. In other words, DNIP with COBiassingle can significantly minimize accuracy differences between odd classes and ground-truth classes, but DNIP with COBias is even more effective when it comes to balancing all pairs of classes, not limited to the odd classes. In summary, optimizing either COBiassingle or COBias is an effective inference-time measure to reduce over-predictions in raw ICL outputs and boost accuracy for under-predicted classes.
4.6 Accuracy-related Objectives vs. the COBias Objective
@@ 4.5 Optimizing COBiassingle
We show that the integration of accuracy-related objectives and the COBias objective are indispensable in achieving the best of both worlds in Figure 5. DNIP objective function $z$ consists of three terms: the error rate term $z_{1}=\frac{1}{M}\sum_{m=1}^{M}\mathbbm{1}\{\hat{y}_{m}\neq y_{m}\}$; the COBias term $z_{2}={N\choose 2}^{-1}\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}|A_{i}-A_{j}|$; and the PMI term $z_{3}=\sum_{j=1}^{N}\textrm{PMI}_{j}$; the accuracy-related terms are $z_{1}$ and $z_{3}$. We ablate all ablation objective functions to perform DNIP: $z_{1},z_{2},z_{3},z_{1}+\beta z_{2},z_{1}-\tau z_{3},\beta z_{2}-\tau z_{3},z_{1}+\beta z_{2}-\tau z_{3}$. Ablations on GPT-2-XL and Llama-2-7B show similar findings (Appendix E). To make fair comparisons, we fix $\beta=2.7,\tau=0.2$ and weight scale $K=30$, which may not result in the same scores reported in Table 3. Among all ablations, $z_{1}+\beta z_{2}-\tau z_{3}$ achieves a balance point between accuracy and COBias. For the criteria, lower COBias is our top priority; when COBias scores are similarly low, we prefer the one that achieves higher accuracy. Compared to $z_{1}+\beta z_{2}-\tau z_{3}$$z_{1}$: (1) Using $z_{1}$ or $z_{3}$ solely emphasizes improving accuracy; though obtaining higher or similar accuracy, they suffer a higher COBias. (2) Using $z_{1}-\tau z_{3}$ also obtains a higher/similar accuracy but a higher COBias. (3) Using $z_{2}$ solely focuses on mitigating COBias and is straightforwardly effective in reducing COBias. However, the shortcoming is obvious: a lower accuracy. (4) Using $z_{1}+\beta z_{2}$ or $\beta z_{2}-\tau z_{3}$ is promising with a similar/lower accuracy and a slightly lower COBias. We note that accuracy-related objectives alone increase accuracy by over 16% ($75\%-59\%$), which is greater than using the objective combining COBias, suggesting DNIP is effective for those who emphasize accuracy improvement.
@@
Test accuracy is improved with only 10 optimization examples, and plateaus at around 500 examples. Figure 7 shows that, at 10 examples, the relative increase in accuracy ranges from 0.2% to 11% except DDI, which is 160%. At 100 examples, the increase ranges from 3% to 208% (DDI, $7.2\%\rightarrow 18.7\%$). The increase continues to grow and stabilizes at around 500 examples. At 1,000 examples, test accuracy for DDI rises to 52%, which is more than 6 times of the ICL accuracy, showing that DNIP is effective for this challenging biomedical NLP task.
@@
Moreover, we observe for RTE a declining trend of accuracy increase, which matches its growing trend of COBias reduction, reflecting the trade-off between the two objectives. We also note a slight decrease in accuracy improvement for SST-5 and DDI at full size, which matches their sudden increase in COBias reduction, verifying that DNIP has emergent COBias reduction abilities with thousands of optimization examples. Overall, DNIP improves both accuracy and COBias at 10 examples, and achieves acceptable COBias reduction and accuracy improvement at 100 examples, with test COBias reduced by 57% and test accuracy improved. For detailed scores, please refer to Appendix F.
@@ 5 Related Work
LLMs may fail in various ways. Analysis towards understanding LLM failure modes has gained remarkable insights on how they fail and how to identify/avoid failures. For example, understanding hallucinations in LLMs facilitates fact-checking evaluation metrics for LLMs.[18, 19, 20]. Jones and Steinhardt [21] reveals failure modes of LLMs which resemble cognitive biases and measures the failures on GPT-3.5 and Codex. Lu et al. [22] identify ICL order sensitivity that result in LLM prompting failures. Turpin et al. [23] shows that LLM reasoning abilities are affected by adding biasing features to the input. Along this line, we tackle a common LLM failure of the class accuracy difference, via nonlinear integer programming optimization on the evaluation metric scores. The inference-time class accuracy bias has multi-faceted causes. It could also be attributed to pre-training data, such as predilections for common tokens in the pretraining corpora [24]. It could be a direct outcome of LLMs’ liking for specific patterns or answers, leading to highly imbalanced predicted class distributions [25]. Other causes include choices of prompt templates [26, 27, 28], demonstrations [29], and prompt orders [22, 23]. Unintended LLM behaviors may also exacerbate the bias, such as user-following sycophancy [30, 31, 32] and prompt toning effects [21, 33, 34, 35].